# SANA 1.5: Efficient Scaling of Training-Time and Inference-Time Compute in Linear Diffusion Transformer

### v2

@@ Abstract
Abstract: This paper presents SANA-1.5, a linear Diffusion Transformer for efficient scaling in text-to-image generation. Building upon SANA-1.0, we introduce three key innovations: (1) Efficient Training Scaling: A depth-growth paradigm that enables scaling from 1.6B to 4.8B parameters with significantly reduced computational resources, combined with a memory-efficient 8-bit optimizer. (2) Model Depth Pruning: A block importance analysis technique for efficient model compression to arbitrary sizes with minimal quality loss. (3) Inference-time Scaling: A repeated sampling strategy that trades computation for model capacity, enabling smaller models to match larger model quality at inference time. Through these strategies, SANA-1.5 achieves a text-image alignment score of 0.72 on GenEval, which can be further improved to 0.80 through inference scaling, establishing a new SoTA on GenEval benchmark. These innovations enable efficient model scaling across different compute budgets while maintaining high quality, making high-quality image generation more accessible. Our code and pre-trained models will be released.
@@ 1 Introduction
Third, we propose an inference-time scaling strategy for SANA, which enables smaller models to match larger model quality through compute rather than parameter scaling. By generating multiple samples and leveraging a VLM-based selection mechanism, our approach improves the GenEval score from 0.72 to 0.80. This improvement follows a similar log-linear scaling pattern observed in LLMs [4], demonstrating that computational resources can be effectively traded for model capacity, challenging the conventional wisdom that larger models are always necessary for better quality.
These three technical contributions - model growth, model depth pruning and inference scaling - form a coherent framework for efficient model scaling. The model growth strategy first explores a larger optimization space, discovering better feature representations. The model depth pruning then identifies and preserves these essential features, enabling efficient deployment. Meanwhile, inference-time scaling provides a complementary perspective. When model capacity is constrained, we can utilize extra inference-time computational resources to achieve similar or even better results than larger models. Together, these approaches demonstrate that thoughtful optimization strategies can outperform simple parameter scaling, providing multiple paths to achieve high quality under different resource constraints.
@@
Our extensive experiments demonstrate that SANA-1.5 achieves 2.5$\times$ faster training convergence than the traditional approach (i.e., scale up and train from scratch). Through our training scaling strategy, we improve the GenEval score from 0.66 to 0.72, which can be further boosted to 0.80 with inference scaling, establishing a new state-of-the-art on the GenEval benchmark. More importantly, our findings reveal a fundamental insight: efficient scaling can be achieved through better optimization trajectories rather than simply increasing model capacity. By leveraging knowledge from smaller models and carefully designing the growth-pruning process, we show that the path to better quailty does not always require larger models.
@@ 2.5 Inference-Time Scaling
To find images that best match a given prompt, we need a model that understands both text and images. While popular models like CLIP [10] and SigLIP [11] offer multi-modal capabilities, their small context windows (77 tokens for CLIP and 66 for SigLIP) limit their effectiveness. This limitation poses a problem since SANA requires long, detailed descriptions. To address this, we explored Visual Language Models to evaluate prompt-matching in generated images. We tested commercial multi-modal APIs, specifically GPT-4o [12] and Gemini-1.5-pro [13], but encountered two significant issues. First, when evaluating single images against prompts, both APIs lacked consistency in their ratings across different runs. Second, when tasked with selecting the best-matching images from multiple options, both models exhibited a strong bias toward the first-presented options, regardless of image ordering or shuffling.
To address these issues, we selected NVILA-2B [14] and developed a specialized dataset to fine-tune it for evaluating images. The fine-tuned model can assess how well images match their prompts while providing detailed explanatory reasoning. The finetuning details are attached in Appendix B. With finetuned VLM to automatically compare and judge generated images, we run a tournament-style comparison several rounds until we determine the top-N candidates, as illustrated in Figure 3(c). This ensures robust selection outcomes and effectively filters out prompt-mismatching images.
@@ 3 Experiments
Method Overall Single Two Counting Colors Position Attribution
@@
SANA-1.5 4.8B 0.72 0.99 0.85 0.77 0.87 0.34 0.54
   + Inference Scaling 0.80 0.99 0.88 0.77 0.90 0.47 0.74
@@ 3.1 Experimental Setup
Evaluation Protocol. We adopt multiple evaluation metrics, including FID, CLIP Score, GenEval [22], and DPG Bench [23], comparing it with SOTA methods. FID and Clip Score are evaluated on the MJHQ-30K [17] dataset, which contains 30K images from Midjourney. GenEval and DPG-Bench both focus on measuring text-image alignment, with 533 and 1,065 test prompts, respectively. We particularly emphasize the GenEval as it better reflects text-image alignment and shows more room for improvement than other metrics.
@@ 3.2 Main Results
We compare SANA-4.8B with the most advanced text-to-image generation methods in Table 1. The scaling from SANA-1.6B to 4.8B brings substantial improvements: 0.06 absolute gains in GenEval (from 0.66 to 0.72), 0.34 reduction in FID (from 5.76 to 5.42), and 0.2 improvement in DPG score (from 84.8 to 85.0). Compared to state-of-the-art methods, our 4.8B model achieves comparable or better results than much larger models like Playground v3 (24B) and FLUX (12B) while using only a fraction of their parameters. Notably, SANA-4.8B demonstrates 0.72 GenEval score, approaching Playground v3’s 0.76, but with 5.5 times lower latency than FLUX-dev (23.0s). Our model also maintains 6.5 times higher throughput than these larger models (compared to FLUX-dev’s 0.04 samples/s), making it more practical for real-world applications. The speed is tested on one A100 GPU with FP16 Precision.
@@
We incorporate inference scaling with the SANA-1.5 4.8B model and compare it against other large image generation models on the GenEval benchmark (Table 2). By selecting samples from 2048 generated images, the inference-scaled model outperforms naive single-image generation by 8% in overall accuracy, with particularly significant improvements in the “Colors”, “Position”, and “Attribution” sub-tasks. Furthermore, equipped with inference scaling, our 4.8B model outperforms Playground v3 (24B) by 4% in overall accuracy. These results demonstrate that trading inference efficiency can enhance model generation quality and accuracy, even with limited model capacity.
@@ 3.3 Analysis
Inference Scaling Law
Figure 8 demonstrates the benefits of scaling up inference-time computation. First, SANA’s accuracy on GenEval consistently improves with more samplings. Second, inference-time scaling enables smaller SANA models to match or even surpass the accuracy of larger ones (1.6B + scaling is better than 4.8B). This reveals the potential of scaling up inference and allows SANA to push toward new state-of-the-art results. As shown in Table 2, our best SANA model with inference scaling outperforms all previous commercial and community models. The only limitation is increased computational cost: sampling $N$ images requires $N\times 49,140G$ FLOPs for SANA generation and $Nlog(N)\times 4,518G$ FLOPs for NVILA judgment and comparison. We leave the efficiency for future work.
@@ Appendix A Full Related Work
Appendix B VLM Finetuning Details for Inference-Time Scaling
To finetune NVILA [14] for SANA inference-time scaling, we generated 2M images and evaluated their alignment with their prompts. We then grouped them by correctness and sampled two candidates from these groups (one positive and one negative). For more stable results, we structured the data in a comparison format and incorporated them into a multimodal conversation format as shown below:
  • User: You are a helpful AI assistant to figure out high-quality AI-generated images. Please pick one that best matches the given prompt: <sana-prompt> <sana-image1> <sana-image2>
  • NVILA-Raw: The first one is better. The latter one generates one extra bag on the left therefore mismatch the prompt.
  • NVILA-JSON: {"pick": 0, "reason": "The latter one generates one extra bag on the left therefore mismatch the prompt"}
Note we alternates the response format between plain text and JSON structure to make the VLM can generate both formats. We totally generate 50k samples and finetune for one epoch. We follow the setting in NVILA with learning $2\times 10^{-5}$, adam optimizer, cosine scheduler with warmup ratio 0.03 and batch size of 8 per device. During inference, we integrate XGrammar [54] to guarantee JSON outputs for comparison.

### v3

@@ Abstract
Abstract: This paper presents SANA-1.5, a linear Diffusion Transformer for efficient scaling in text-to-image generation. Building upon SANA-1.0, we introduce three key innovations: (1) Efficient Training Scaling: A depth-growth paradigm that enables scaling from 1.6B to 4.8B parameters with significantly reduced computational resources, combined with a memory-efficient 8-bit optimizer. (2) Model Depth Pruning: A block importance analysis technique for efficient model compression to arbitrary sizes with minimal quality loss. (3) Inference-time Scaling: A repeated sampling strategy that trades computation for model capacity, enabling smaller models to match larger model quality at inference time. Through these strategies, SANA-1.5 achieves a text-image alignment score of 0.81 on GenEval, which can be further improved to 0.96 through inference scaling with VILA-Judge, establishing a new SoTA on GenEval benchmark. These innovations enable efficient model scaling across different compute budgets while maintaining high quality, making high-quality image generation more accessible. Our code and pre-trained models are released.
@@ 1 Introduction
Third, we propose an inference-time scaling strategy for SANA, which enables smaller models to match larger model quality through compute rather than parameter scaling. By generating multiple samples and leveraging a VLM-based selection mechanism, our approach improves the GenEval score from 0.81 to 0.96. This improvement follows a similar log-linear scaling pattern observed in LLMs [4], demonstrating that computational resources can be effectively traded for model capacity, challenging the conventional wisdom that larger models are always necessary for better quality.
rThese three technical contributions - model growth, model depth pruning and inference scaling - form a coherent framework for efficient model scaling. The model growth strategy first explores a larger optimization space, discovering better feature representations. The model depth pruning then identifies and preserves these essential features, enabling efficient deployment. Meanwhile, inference-time scaling provides a complementary perspective. When model capacity is constrained, we can utilize extra inference-time computational resources to achieve similar or even better results than larger models. Together, these approaches demonstrate that thoughtful optimization strategies can outperform simple parameter scaling, providing multiple paths to achieve high quality under different resource constraints.
@@
Our extensive experiments demonstrate that SANA-1.5 achieves 2.5$\times$ faster training convergence than the traditional approach (i.e., scale up and train from scratch). Through our training scaling strategy, we improve the GenEval score from 0.66 to 0.81, which can be further boosted to 0.96 with inference scaling, establishing a new state-of-the-art on the GenEval benchmark. More importantly, our findings reveal a fundamental insight: efficient scaling can be achieved through better optimization trajectories rather than simply increasing model capacity. By leveraging knowledge from smaller models and carefully designing the growth-pruning process, we show that the path to better quailty does not always require larger models.
@@ 2.5 Inference-Time Scaling
To find images that best match a given prompt, a model that understands both text and images is needed. While popular models like CLIP [10] and SigLIP [11] offer multi-modal capabilities, their small context windows (77 tokens for CLIP and 66 tokens for SigLIP) limit their effectiveness. This limitation poses a problem since SANA usually takes long, detailed descriptions as inputs. To address this, we explored Visual Language Models to evaluate prompt-matching for generated images. We tested commercial multi-modal APIs, specifically GPT-4o [12] and Gemini-1.5-pro [13], but encountered two significant issues. First, when evaluating single images against prompts, both APIs lacked consistency in their ratings across different runs. Second, when tasked with selecting the best-matching images from multiple options, both models exhibited a strong bias toward the first-presented options, regardless of image ordering or shuffling.
Instead of applying existing models or APIs, we trained a specialized NVILA-2B [14] to score the images, which we named VILA-Judge: Sana generates then VILA picks. We created a 2M prompt-matching dataset with prompts generated in the GenEval style. We excluded the prompts that are already existed in GenEval evalset and generated the images using using Flux-Schnell [15] to avoid overfitting. We then formatted these as multimodal conversations, as shown below
  • User: You are an AI assistant specializing in image analysis and ranking. Your task is to analyze and compare image based on how well they match the given prompt. <image> The given prompt is: <prompt>. Please consider the prompt and the image to make a decision and response directly with ’yes’ or ’no’
  • VILA-Judge: ’yes’ / ’no’.
The fine-tuned VILA-Judge can effectively assess how well images match their prompts and robustly filters out prompt-mismatching images. During inference, we compare two images in each round until the top-N candidates are determined
  • When VILA-Judge responses one ‘yes’ and one ‘no’, we pick the image with ‘yes’;
  • When VILA-Judge responses both ‘yes’ or ‘no’, we pick the image with higher confidence (logprob);
Such tournament-style comparison robustly filters out prompt-mismatching images and consistently boosts the GenEval scores, as illustrated in Figure 3(c) and Table. 2. More details of inference time scaling are attached in Appendix B.
@@ 3 Experiments
Method Overall Single Two Counting Colors Position Color Attribution
@@
SD1.5 [22] 0.42 0.98 0.39 0.31 0.72 0.04 0.06
   + Inference Scaling 0.87 1.00 0.97 0.93 0.96 0.75 0.62
SANA-1.5 4.8B Pre 0.72 0.99 0.85 0.77 0.87 0.34 0.54
SANA-1.5 4.8B Ours 0.81 0.99 0.93 0.86 0.84 0.59 0.65
   + Inference Scaling 0.96 1.00 1.00 0.97 0.94 0.96 0.87
@@ 3.1 Experimental Setup
Evaluation Protocol. We adopt multiple evaluation metrics, including FID, CLIP Score, GenEval [23], and DPG Bench [24], comparing it with SOTA methods. FID and Clip Score are evaluated on the MJHQ-30K [18] dataset, which contains 30K images from Midjourney. GenEval and DPG-Bench both focus on measuring text-image alignment, with 553 and 1,065 test prompts, respectively. We particularly emphasize the GenEval as it better reflects text-image alignment and shows more room for improvement than other metrics.
@@ 3.2 Main Results
We compare SANA-4.8B with the most advanced text-to-image generation methods in Table 1. The scaling from SANA-1.6B to 4.8B Pre (Pre-trained) brings substantial improvements: 0.06 absolute gains in GenEval (from 0.66 to 0.72), 0.34 reduction in FID (from 5.76 to 5.42), and 0.2 improvement in DPG score (from 84.8 to 85.0). Compared to state-of-the-art methods, our 4.8B model achieves comparable or better results than much larger models like Playground v3 (24B) and FLUX (12B) while using only a fraction of their parameters. We further perform post-training on the 4.8B model using a high-quality dataset, with details provided in Sec.  3.3. To differentiate between the models, we refer to the pre-trained model as Pre and the post-trained model as Ours. After post-training, SANA-4.8B achieves a significant performance improvement on the GenEval benchmark, increasing from 0.72 to 0.81. Notably, SANA-4.8B demonstrates 0.81 GenEval score, outperforming Playground v3’s 0.76, but with 5.5 times lower latency than FLUX-dev (23.0s). Our model also maintains 6.5 times higher throughput than these larger models (compared to FLUX-dev’s 0.04 samples/s), making it more practical for real-world applications. The speed is tested on one A100 GPU with FP16 Precision.
@@
We incorporate inference scaling with the SANA-1.5 4.8B v2 model and compare it against other large image generation models on the GenEval benchmark (Table 2). By selecting samples from 2048 generated images, the inference-scaled model outperforms naive single-image generation by 15% in overall accuracy from 0.81 to 0.96, with particularly significant improvements in the “Position” (from 0.59 to 0.96), and “Color Attribution” (from 0.65 to 0.87). Furthermore, equipped with inference scaling, our 4.8B model outperforms Playground v3 (24B) by 20% in overall accuracy (0.76 vs 0.96). These results demonstrate that trading inference efficiency can enhance model generation quality and accuracy, even with much smaller base model size (4.8B vs 24B).
@@ 3.3 Analysis
Furthermore, we created a new small-scale SFT data set, consisting of 144,291 images generated from 18,240 prompts in the GenEval style. Each prompt generates 7.91 “correct” images on average. We limit each prompt to generate at most 10 correct images to prevent overfitting. The correct image is filtered by the GenEval toolkit. Note that we excluded the prompts already existing in GenEval evalset and generated images using Flux-Schnell to avoid overfitting. As shown in Table 2, this new SFT data improves the GenEval result of SANA-4.8B from 0.72 (v1) to 0.81 (v2).
Inference Scaling Law
Figure 8 demonstrates the benefits of scaling up inference-time computation. First, SANA’s accuracy on GenEval consistently improves with more samplings. Second, inference-time scaling enables smaller SANA models to match or even surpass the accuracy of larger ones (1.6B + scaling is better than 4.8B). This reveals the potential of scaling up inference and allows SANA to push toward new state-of-the-art results. As shown in Table 2, our best SANA model with inference scaling outperforms all previous commercial and community models. The only limitation is increased computational cost: sampling $N$ images requires $N\times 49,140G$ FLOPs for SANA generation and $2N\times 4,518G$ FLOPs for VILA-Judge judgment and comparison. We leave the efficiency for future work.
@@ Appendix A Full Related Work
Appendix B Inference-Time Scaling Details
Dataset. To finetune NVILA [14] for SANA inference-time scaling, we generated 2.5M images and evaluated their alignment with the given prompts using the GenEval toolkit. We utilize 15,654 unique prompts and generate 160 images per prompt using Flux-Schnell. Our prompts are constructed in a style similar to GenEval, incorporating 80 object classes from Mask2Former along with attributes such as color, spatial location, quantity, and various relationships. Importantly, our prompts do not overlap with the GenEval test set. To evaluate the generated images, we leverage the GenEval toolkit. Finally, we train NVILA using the prompts, images, and their corresponding labels (yes or no).
Training Setup. We follow the setting in NVILA with learning $2\times 10^{-5}$, Adam optimizer, cosine scheduler with warmup ratio 0.03 and batch size of 8 per device and train the 2M dataset with one epoch.